# Investigating the role of nuclear parameters on oscillation modes in hot Neutron Stars

### v1

@@ I Introduction
Astrophysical events such as proto-neutron stars (PNSs) born out of supernovae (SN) and binary mergers exceed temperatures of tens of MeV and the cold matter approximations are no longer valid  [29, 30, 31, 32, 33]. It is important to take into account thermal effects as they play a significant role in supernova dynamics, oscillations in PNS as well as in determining stability of the post-merger remnants, which along with the composition can change the threshold mass limit for prompt collapse to a black hole [34]. The EoS of a finite temperature system can be uniquely characterized by three quantities, temperature $T$, baryon number density $n_{b}$ and the charge fraction of strongly interacting particles $Y_{Q}$. In order to be applicable in numerical simulations, a wide range and finely spaced sets of parameters should be considered, along with different possible composition in the neutron star interior. The EoS of hot homogeneous matter spans 0-100 MeV in temperature, 0.01-0.60 in charge fraction and 0.5-6.0 times in nuclear saturation densities in BNS mergers and supernovae [6]. In the absence of EoS models spanning the required parameter ranges and composition, the so-called $\Gamma$-law (supplementing the cold EoS with ideal gas thermal contributions) has often been employed to incorporate finite temperature effects in simulations. However, this method of constructing EoS has been shown to be inappropriate particularly for matter consisting of non-nucleonic degrees of freedom [35, 36, 37, 38]. It has been described in a recent work in quark-hadron phase transition [38] that $\Gamma$ value should be different in hadronic ($\Gamma\sim 1.75$) and quark ($\Gamma\sim 4/3$) phases. The $\Gamma$-law also does not take into account the effect of charge fraction across all densities. Therefore it is crucial to incorporate composition and thermal EoS effects consistently into simulations.
@@
Cold NSs are known to follow certain EoS independent universality such as $C$-Love, $\bar{I}-\Lambda-\bar{Q}$, $f$-Love [67, 68, 49] among others. Much effort has been dedicated into deriving universal relations (URs) that link the oscillation mode properties of PNSs with their physical characteristics, such as mass and radius, which are crucial for GW asteroseismology [58, 56, 61] of proto-neutron stars. Raduta et al. [69] obtained $\bar{I}-\Lambda-\bar{Q}-C$ relations for hot NSs using a DD-RMF parameterization. This was extended by Khadkikar et al. [70] to explore URs between static and Kepler rotation properties such as mass, radius and frequency etc. and fit relations were provided for $\bar{I}-C$ and $\bar{Q}-C$ for different thermal configurations with various tabulated EoSs. Similar UR study for nucleonic and hybrid hot stars were performed by Largani et al. [71] where they provided fits for normalized moment of inertia with compactness at Keplerian limit for fixed values of entropy per baryon and lepton fraction. Laskos-Patkos et al. [72] studied deviations of hot neutron stars from cold universality for $C$-Love relations.
@@ II.1 Equation of State
where $\mu_{i}$ is the actual chemical potential of the corresponding species and can be written in terms of baryon ($\mu_{B}$) and charge ($\mu_{Q}$) chemical potentials. If $B_{i}$ and $Q_{i}$ are baryon and electric charges of the $i$-th species then, $\mu_{i}=B_{i}\mu_{B}+Q_{i}\mu_{Q}$. Both particle and anti-particle species contributes to entropy, as do leptons/anti-leptons and radiation. The effective chemical potentials for anti-particles are negative of corresponding particle effective chemical potentials. Chemical potential for electrons is $\mu_{e}=\mu_{L}-\mu_{Q}$, where $\mu_{L}$ is the lepton chemical potential. The total entropy density ($s$) is given by
@@
The direct consequences of finite temperature are effects such as out-of-$\beta$ equilibrium, neutrino trapping among others. These effects can manifest themselves in the hot EoS and consequently in the macroscopic properties. At postbounce evolution in PNSs, the thermal properties control the EoS until the neutron star becomes cold and degenerate. These thermal properties are entropy per baryon ($S/A$) and electric charge (out-of-$\beta$ equilibrium) and/or lepton fractions (in case of neutrino trapping). Initially the star is in a low entropy state with high charge/lepton fractions, it then becomes a hot (high entropy) star with low charge/lepton fractions and eventually attains a deleptonized cold $\beta$ equilibrium state [51]. It has been shown in previous works that this leads to deviations from universal relations (URs) usually observed in cold neutron stars [69, 77, 46].
The core temperature profiles corresponding to different thermodynamic conditions for nucleonic matter $N(S/A,Y_{Q})$ i.e., different values of entropy per baryon $S/A$ and charge fraction $Y_{Q}$, are shown in fig. 1. It can be seen that higher entropy leads to larger temperatures and higher proton fraction results in lower temperatures at high densities.
@@ II.3 Oscillation modes
We display $f$- and $p_{1}$-mode frequencies of some specific thermal cases in figs. 3(a) and 3(b) respectively calculated in general relativistic treatment. The $p_{1}$-mode frequencies are highest for the cold configuration and they are found to decrease with an increase in entropy per baryon and charge fraction. The isothermal configuration for $T=20$ MeV has the lowest $p_{1}$-mode frequencies. A crossing appears between $N(1,0.4)$ and $N(2,0.2)$ configurations for the $p_{1}$-mode at high mass. This behaviour is also seen in the full parameter space allowed by various constraints (discussed in subsection III.3). On the other hand, $f$-mode frequencies for all cases have some overlapping regions at intermediate and high mass ranges.
@@ II.4 Constraining the model parameters
Table 1: Uncertainty ranges in nuclear saturation parameters used in this study: the 2nd row provides the lower limit and 3rd row gives upper limit [15].
@@ III Results
The composition as well as thermodynamical properties of the neutron star change with the evolving temperature during the PNS formation or neutron star merger. To investigate the effects of temperature and out-of-$\beta$ equilibrium, in this work we consider two thermal configurations, $N(S/A=1,Y_{Q}=0.4)$ and $N(S/A=2,Y_{Q}=0.2)$ representing two different stages of evolution of a hot PNS. The earlier evolution state corresponds to a hot lepton/electron rich configuration while at later times the state corresponds to a hotter core with deleptonized configuration [94, 29, 95]. We also show comparison with cold $\beta$ equilibrated state in our study.
@@ III.4.1 $C$-Love Relations
Figure 16: $\Lambda$ v/s $C$ universality for the three cases.
@@
Figure 17: Effect of $S/A$ on $\Lambda-C$ universality. Nuclear saturation parameters are fixed to same values as in fig. 1. $S/A$ has been varied between $1-$ in steps of 0.01.
It is clear that $\Lambda-C$ relation depends not only on charge fraction but also on whether the matter is in $\beta$ equilibrium when we vary $S/A$. The spread in universality in fig. 17 is much less than that of fig. 18. From the results in fig. 17 one can obtain universal fits similar to the works of Maselli et al. [67]. We provide the universal fits for the two thermal configurations in Table 2. On the other hand, we see from fig. 18 that $Y_{Q}$, which is a measure of out-of-$\beta$ equilibrium, is mainly responsible for breaking this universality and there is no possibility to obtain a fit relation.
@@
These findings are reasonable, as at higher $\Lambda$ the star can be assumed to be in the $N(1,0.4)$ representative state and this configuration shows the highest deviation from cold $f$-Love universality. The $N(2,0.2)$ state is representative of intermediate $\Lambda$ and it has close resemblance with the cold $\beta$ equilibrated configuration. The close resemblance these URs have with $N(T=0)$ promotes us to study them across ranges of $S/A$ and $Y_{Q}$. The conclusions are somewhat different to those for $C-Love$ relations (see Appendix subsections V.1 and V.2).
@@ IV Discussions
Previous studies have indicated deviations of $\bar{I}-\bar{Q}$ relations in hot environment [46, 69, 107, 108]. A few studies [69, 70] showed that universality is restored if finite $T$ configurations for different EoS with same $S/A$ and electron fraction are considered. PNS simulations by Sotani et al. [109] explored the prospect of measurement PNS mass and radius using simultaneous measurements of $f$- (or $p_{1}$-) and $w_{1}$-mode frequencies by relating those to average mass density and stellar compactness respectively. Torres-Forné et al. [58] have demonstrated how $g_{1}$ and $f$-mode frequencies scale with surface gravity in their PNS simulations. In our article, we have shown that $C$-Love and $f$-Love URs are mostly not sensitive to nuclear saturation properties but they change depending on the thermal state. In particular charge fraction significantly affects these deviations and they occur at high $\Lambda$-values (or low compactness). A recent simulation [110] has also shown that due to thermal effects deviations from URs i.e., $f$-$\sqrt{M/R^{3}}$, $\bar{I}-\Lambda-\bar{Q}$, $f$-Love occur during PNS evolution at earlier stages where the configuration has large tidal deformability.
@@
This study can be extended to include non-nucleonic degrees of freedom such as hyperons, delta baryons and quark matter in hot EoS that satisfy microscopic, experimental and astrophysical constraints. These new degrees of freedom are known to soften the EoSs and might also lead to decrease in core temperature due to entropy being contributed by these new species. It would also be interesting to include the effect of neutrinos as they are expected to be in equilibrium with the thermal/isentropic bath at high temperatures.
V appendix
V.1 Effect of entropy per baryon on $f$-Love relations
@@
We fix the nuclear saturation properties as in fig. 1. Figs. 24 and 25 show effect of changing $S/A$ between $1-2$ in $Y_{Q}=0.4$ and $\beta$ equilibrated states. From fig. 24, we see that $Y_{Q}=0.4$ case shows deviations from $\beta$ equilibtrated hot NSs for higher $\Lambda$ and they are overlapping otherwise. This is in contrast to $\Lambda-C$ URs shown in fig. 17 where we did not get any overlapping region. The relation between $\rm{Im}$ ($M\omega)$ and $\ln(\Lambda)$ shown in fig. 25 also establishes deviations at high $\Lambda$ values. There are also qualitatively negligible deviations for both real and imaginary parts across the entropy per baryon range considered as compared to $\Lambda-C$ URs.
@@
These deviations shown between states with $\beta$ equilibration and $Y_{Q}=0.4$ clearly shows that at higher $\Lambda$ universality is broken by out-of-$\beta$ equilibrium effects. The spreads however in figs. 24 and 25 are qualitatively much lower than that of $\Lambda-C$ relations shown in fig. 18 where deviations are observed across the entire range of dimensionless tidal deformability.
Figure 24: Effect of $S/A$ on $\rm{Real}$ ($M\omega_{f})-\Lambda$ universality. Nuclear saturation properties are same as in fig. 1. $S/A$ has been varied between 1-2 in steps of 0.01.
Figure 25: Effect of $S/A$ on $\rm{Im}$ ($M\omega_{f})-\Lambda$ universality. Nuclear saturation properties are same as in fig. 1. $S/A$ has been varied between $1-2$ in steps of 0.01.
V.2 Effect of charge fraction on $f$-Love relations
@@
Here we fix entropy per baryon at $S/A=1$ and vary $Y_{Q}$ from 0.01 - 0.40. We see in figs. 26(a) and 26(b) that this has no effect on both real and imaginary parts of $f$-Love relations for low $\Lambda$ values and at higher values there are small spreads. These deviations appear for $\Lambda>e^{8}\sim 3000$. The conclusion is that charge fraction is responsible for deviations in the universality like $C$-Love relations albeit to a much lesser degree and only at higher $\Lambda$.
@@
Acknowledgements
The authors are grateful to Suprovo Ghosh for providing the posterior data that was used in this work. D.C. would like to thank Adriana R. Raduta and Mikhail V. Beznogov for their warm hospitality at IFIN-HH, Bucharest which led to several interesting discussions related to this work. D.C. would also like to thank Prof. J. Schaffner-Bielich and his students at ITP Frankfurt for insightful discussions on finite temperature models of neutron stars. N.B., B.K.P. and D.C. acknowledge the usage of the IUCAA HPC computing facility for numerical calculations of oscillation modes.

### v2

@@ I Introduction
Astrophysical events such as proto-neutron stars (PNSs) [29, 30, 31, 32] born out of supernovae (SN) and binary neutron star (BNS) mergers [33, 34, 35, 36, 37, 38] exceed temperatures of tens of MeV and the cold matter approximations are no longer valid. It is important to take into account thermal effects as they play a significant role in supernova dynamics, oscillations in PNS as well as in determining stability of the post-merger remnants, which along with the composition can change the threshold mass limit for prompt collapse to a black hole [39]. The EoS of a finite temperature system can be uniquely characterized by three quantities, temperature $T$, baryon number density $n_{b}$ and the charge fraction of strongly interacting particles $Y_{Q}$. In order to be applicable in numerical simulations, a wide range and finely spaced sets of parameters should be considered, along with different possible composition in the neutron star interior. The EoS of hot homogeneous matter spans 0-100 MeV in temperature, 0.01-0.60 in charge fraction and 0.5-6.0 times in nuclear saturation densities in BNS mergers and supernovae [6]. In the absence of EoS models spanning the required parameter ranges and composition, the so-called $\Gamma$-law (supplementing the cold EoS with ideal gas thermal contributions) has often been employed to incorporate finite temperature effects in simulations. However, this method of constructing EoS has been shown to be inappropriate particularly for matter consisting of non-nucleonic degrees of freedom [40, 41, 42, 43]. It has been described in a recent work in quark-hadron phase transition [43] that $\Gamma$ value should be different in hadronic ($\Gamma\sim 1.75$) and quark ($\Gamma\sim 4/3$) phases. The $\Gamma$-law also does not take into account the effect of charge fraction across all densities. Therefore it is crucial to incorporate composition and thermal EoS effects consistently into simulations.
@@
Cold NSs are known to follow certain EoS independent universality such as $C$-Love, $\bar{I}-\Lambda-\bar{Q}$, $f$-Love [72, 73, 54, 74, 75, 76] among others. Much effort has been dedicated into deriving universal relations (URs) that link the oscillation mode properties of PNSs with their physical characteristics, such as mass and radius, which are crucial for GW asteroseismology [63, 61, 66, 77] of proto-neutron stars. Raduta et al. [78] obtained $\bar{I}-\Lambda-\bar{Q}-C$ relations for hot NSs using a DD-RMF parameterization. This was extended by Khadkikar et al. [79] to explore URs between static and Kepler rotation properties such as mass, radius and frequency etc. and fit relations were provided for $\bar{I}-C$ and $\bar{Q}-C$ for different thermal configurations with various tabulated EoSs. Similar UR study for nucleonic and hybrid hot stars were performed by Largani et al. [80] where they provided fits for normalized moment of inertia with compactness at Keplerian limit for fixed values of entropy per baryon and lepton fraction. Laskos-Patkos et al. [81] studied deviations of hot neutron stars from cold universality for $C$-Love relations.
@@ II.1 Equation of State
where $\mu_{i}$ is the actual chemical potential of the corresponding species and can be written in terms of baryon ($\mu_{B}$) and charge ($\mu_{Q}$) chemical potentials. If $B_{i}$ and $Q_{i}$ are baryon and electric charges of the $i$-th species then, $\mu_{i}=B_{i}\mu_{B}+Q_{i}\mu_{Q}$. Both particle and anti-particle species contributes to entropy, as do leptons/anti-leptons and radiation. The effective chemical potentials for anti-particles are negative of corresponding particle effective chemical potentials. Chemical potential for electrons is $\mu_{e}=\mu_{L}-\mu_{Q}$, where $\mu_{L}$ is the lepton chemical potential. The total entropy density ($s$) is given by,
@@
The direct consequences of finite temperature are effects such as out-of-$\beta$ equilibrium, neutrino trapping among others. These effects can manifest themselves in the hot EoS and consequently in the macroscopic properties. At postbounce evolution in PNSs, the thermal properties control the EoS until the neutron star becomes cold and degenerate. These thermal properties are entropy per baryon ($S/A$) and electric charge (out-of-$\beta$ equilibrium) and/or lepton fractions (in case of neutrino trapping). During PNS evolution and also BNS merger scenarios, the system goes through different stages of $S/A$ and $Y_{Q}$. We discuss two such thermal configurations relevant for PNS evolution in section III and show the effect of variation of nuclear properties on these configurations subsequently. It has been shown in previous works that this leads to deviations from universal relations (URs) usually observed in cold neutron stars [78, 86, 51].
To verify expected behaviour of thermal properties as given in existing literature, we first use a fixed nuclear parameterization (from table 1) and vary only the thermal properties of the star. The core temperature profiles corresponding to different thermodynamic conditions for nucleonic matter $N(S/A,Y_{Q})$ i.e., different values of entropy per baryon $S/A$ and charge fraction $Y_{Q}$, are shown in fig. 1. It can be seen that higher entropy leads to larger temperatures and higher proton fraction results in lower temperatures at high densities.
@@ II.3 Oscillation modes
We display $f$- and $p_{1}$-mode frequencies of some specific thermal cases in Fig. 3(a) and 3(b) respectively calculated in general relativistic treatment. The $p_{1}$-mode frequencies are highest for the cold configuration and they are found to decrease with an increase in entropy per baryon and charge fraction. The isothermal configuration for $T=20$ MeV has the lowest $p_{1}$-mode frequencies. A crossing appears between $N(1,0.4)$ and $N(2,0.2)$ configurations for the $p_{1}$-mode at high mass. This behavior is also seen in the full parameter space allowed by various constraints (discussed in subsection III.3). On the other hand, $f$-mode frequencies for all cases have some overlapping regions at intermediate and high mass ranges. From  fig. 3, it can be concluded that both the $f$- and $p_{1}$-mode frequencies decrease in thermally configured stars compared to cold NSs. Notably, the thermal effects have a more pronounced impact on the $p_{1}$-mode frequencies than on the $f$-mode frequencies. The reduction in mode frequencies with thermal effects aligns with the findings of previous studies  [56, 95, 96]. This behavior can be attributed to the fact that both $f$-mode and $p$-mode frequencies scale inversely with a power of the stellar radius (see [97, 61]). As thermally configured stars of a given mass exhibit lower compactness than their cold counterparts, the increase in radius leads to a reduction in mode frequencies. Thus, the observed variations in frequency are a direct consequence of the underlying differences in compactness between hot and cold configurations (see fig. 2).
@@ II.4 Constraining the model parameters
Table 1: Nuclear saturation parameters considered in this work [15]: The top row is the fixed set of the nuclear parameters used in figures 12 and 3 and the bottom row shows the range of varied parameters used in the rest of this study.
@@ III Results
The composition as well as thermodynamical properties of the neutron star change with the evolving temperature during the PNS formation or neutron star merger. In PNS early evolution the star goes through different stages in which the core evolves from a low temperature (low entropy state) and gets heated up (high entropy) with a decrease in charge fraction [56, 106, 29]. It then starts cooling and eventually becomes cold and attains $\beta$ equilibrium. In case of BNS mergers a recent simulation has shown that the entropy per baryon can reach upto values of 3 [34]. To investigate the effects of temperature and out-of-$\beta$ equilibrium, in this work we consider two thermal configurations, $N(S/A=1,Y_{Q}=0.4)$ and $N(S/A=2,Y_{Q}=0.2)$ representing two different stages of evolution of a hot PNS. The former state corresponds to a hot lepton/electron rich configuration while the latter corresponds to a hotter core with deleptonized configuration [56]. We also show comparison with cold $\beta$ equilibrated state in our study.
@@ III.4.1 $C$-Love Relations
Figure 16: $\Lambda~{}v/s~{}C$ universality for the three cases.
@@
Figure 17: Effect of $S/A$ on $\Lambda-C$ universality. Nuclear saturation parameters are fixed to same values as in fig. 1. $S/A$ has been varied between $1-2$ in steps of 0.01.
It is clear that $\Lambda-C$ relation depends not only on charge fraction but also on whether the matter is in $\beta$ equilibrium when we vary $S/A$. The spread in universality in fig. 17 is much less than that of fig. 18. From the results in fig. 17 one can obtain universal fits similar to the previous works  [72, 75]. We provide the universal fits for the two thermal configurations in Table 2. On the other hand, we see from fig. 18 that $Y_{Q}$, which is a measure of out-of-$\beta$ equilibrium, is mainly responsible for breaking this universality and there is no possibility to obtain a fit relation.
@@ III.4.2 $f$-Love Relations
These findings are reasonable, as at higher $\Lambda$ the star can be assumed to be in the $N(1,0.4)$ representative state and this configuration shows the highest deviation from cold $f$-Love universality. The $N(2,0.2)$ state is representative of intermediate $\Lambda$ and it has close resemblance with the cold $\beta$ equilibrated configuration. The close resemblance these URs have with $N(T=0)$ promotes us to study them across ranges of $S/A$ and $Y_{Q}$. The conclusions are somewhat different to those for $C$-Love relations (see Appendix sections A and B).
@@ IV Discussions
Previous studies have indicated deviations of $\bar{I}-\bar{Q}$ relations in hot environment [51, 78, 116, 117]. A few studies [78, 79] showed that universality is restored if finite $T$ configurations for different EoS with same $S/A$ and electron fraction are considered. PNS simulations by Sotani et al. [97] explored the prospect of measurement PNS mass and radius using simultaneous measurements of $f$- (or $p_{1}$-) and $w_{1}$-mode frequencies by relating those to average mass density and stellar compactness respectively. Torres-Forné et al. [63] have demonstrated how $g_{1}$ and $f$-mode frequencies scale with surface gravity in their PNS simulations. In our article, we have shown that $C$-Love and $f$-Love URs are mostly not sensitive to nuclear saturation properties but they change depending on the thermal state. In particular charge fraction significantly affects these deviations and they occur at high $\Lambda$-values (or low compactness). A recent simulation [118] has also shown that due to thermal effects deviations from URs i.e., $f$-$\sqrt{M/R^{3}}$, $\bar{I}-\Lambda-\bar{Q}$, $f$-Love occur during PNS evolution at earlier stages where the configuration has large tidal deformability. It should be noted that in the case of proto-neutron stars, in addition to thermal effects one must also take into account the effect of strong magnetic fields which are known to break the universality of $\bar{I}-$Love $-\bar{Q}$ relations [119].
@@
This study can be extended to include non-nucleonic degrees of freedom such as hyperons, delta baryons and quark matter in hot EoS that satisfy microscopic, experimental and astrophysical constraints. These new degrees of freedom are known to soften the EoSs and might also lead to decrease in core temperature due to entropy being contributed by these new species.
It would also be interesting to include the effect of neutrinos as they are expected to be in equilibrium with the thermal/isentropic bath at high temperatures. Previous studies on BNS mergers [121, 122] have shown that in the temperature regime $T\sim 1-5$ MeV, nuclear matter is transparent to neutrinos due to their long mean free path. At high temperatures (beyond $\sim 5$ MeV) neutrinos may be trapped, depending on the neutrino mean free path at a given density. As a first step, in this work we developed the formalism for the $\nu$-transparent case (valid for temperatures upto $5$ MeV) and plan to extend this to include trapped neutrinos in a future work.
Acknowledgements.
The authors are grateful to Suprovo Ghosh for providing the posterior data that was used in this work. D.C. would like to thank Adriana R. Raduta and Mikhail V. Beznogov for their warm hospitality at IFIN-HH, Bucharest which led to several interesting discussions related to this work. D.C. would also like to thank Prof. J. Schaffner-Bielich and his students at ITP Frankfurt for insightful discussions on finite temperature models of neutron stars. D.C. is grateful to Prof Mark Alford, Alexander Haber and their group at Washington University in St. Louis for their hospitality and fruitful exchanges. The authors also thank Prof Luciano Rezzolla for pointing them to his relevant works. N.B., B.K.P. and D.C. acknowledge the usage of the IUCAA HPC computing facility for numerical calculations of oscillation modes.
V Data availability
The posterior data of both the nuclear saturation properties and the coupling constants obtained after imposing $\chi$EFT + Astro and $\chi$EFT + Astro + HIC constraints which support this work can be found at [123]. Alternatively the posterior sets of only nuclear saturation properties are available in the arXiv ancillary files of previously published work of Ghosh et al. [14].
Appendix A Effect of entropy per baryon on $f$-Love relations
@@
We fix the nuclear saturation properties as in fig. 1. Figs. 24(a) and 24(b) show effect of changing $S/A$ between $1-2$ in $Y_{Q}=0.4$ and $\beta$ equilibrated states. From fig. 24(a), we see that $Y_{Q}=0.4$ case shows deviations from $\beta$ equilibtrated hot NSs for higher $\Lambda$ and they are overlapping otherwise. This is in contrast to $\Lambda-C$ URs shown in fig. 17 where we did not get any overlapping region. The relation between $\rm{Im}$($M\omega)$ and $\ln(\Lambda)$ are shown in fig. 24(b) also establishes deviations at high $\Lambda$ values. There are also qualitatively negligible deviations for both real and imaginary parts across the entropy per baryon range considered as compared to $\Lambda-C$ URs.
[…]
Figure 24: Effect of entropy per baryon on $f$-Love universality.
@@
These deviations shown between states with $\beta$ equilibration and $Y_{Q}=0.4$ clearly shows that at higher $\Lambda$ universality is broken by out-of-$\beta$ equilibrium effects. The spreads however in figs. 24(a) and 24(b) are qualitatively much lower than that of $\Lambda-C$ relations shown in fig. 18 where deviations are observed across the entire range of dimensionless tidal deformability.
Appendix B Effect of charge fraction on $f$-Love relations
@@
Here we fix entropy per baryon at $S/A=1$ and vary $Y_{Q}$ from 0.01 - 0.40. We see in figs. 25(a) and 25(b) that this has no effect on both real and imaginary parts of $f$-Love relations for low $\Lambda$ values and at higher values there are small spreads. These deviations appear for $\Lambda>e^{8}\sim 3000$. The conclusion is that charge fraction is responsible for deviations in the universality like $C$-Love relations albeit to a much lesser degree and only at higher $\Lambda$.